A gardener will plant 3 yellow tulips, 2 red tulips, and 4 purple tulips along a
straight path. How many different arrangements can she make?

Found 2 solutions by HasanSahin, sudhanshu_kmr:
Answer by HasanSahin(52)

You will do factorial in this question..
There 9 places as the sum of the number of tulips are 9.
You'll do 3 different calculations for 3 different colour..
(number of places)! / [(number of tulips in same colour)! * (number of places - number of tulips in same colour)!]
Start with red ones..
9! / [2! * (9-2)!] = a is the number of possibilities for red ones..
Continue with yellows but now you have 9-2=7 places left as you used 2 of them.
7! / [3! * (7-3)!] = b is the number of possibilities for yellow ones..
Continue with purples but now you have 7-3=4 places left as you used 5 of them.
4! / [4! * (4-4)!] = c is the number of possibilities for purple ones..
The total possibility in this question is the the multiplication of a*b*c
RF.

Answer by sudhanshu_kmr(1152)


total no. of tulips = 9

total no. of arrangements = (9!)/[3! * 2! * 4!]


answer is : 1260




It is possible that some typing mistake in solution of a problem, if any please ignore it. Understand the concept and try to solve the problem yourself. If there is problem related to concept, contact at
sudhanshu.cochin@yahoo.com or sudhanshu.cochin@gmail.com
Best of luck.......